Please help me solve this equation:
Deanna want to find the distance across the lake from point B to point C. To do this, she took measurements from point A to both points. The disatance from A to B was 170 yards. The distance from A to C was 80 yards. What was the distance from B to C?

Found 2 solutions by jorel1380, richard1234:
Answer by jorel1380(3719)

(A to B)- (A to C)=170-80=90 yards..

Answer by richard1234(7193)

Do A, B, and C lie on a triangle? Or are they collinear?